2. Find the solution of the second order differential equations: d2 +y = 0, y(T/3) = 0, y(T/3) = 4 a. dx2 b. Y - 8y + 16y
Answer #1

2) a) Given QE do W ty dax? - (071)4 - ib Auxiliary equation is m²71=0 mo ti Har Y = 4 cosx + y sina y=4 G cosa tesina put Y-Beta2 2 2 -Butle=8 kle get solve ☺ & Bei -13 + 34₂ = o 8 t C 46₂ C2 = 2 9 = -3.2 - 213 kle get substitute 9, ce value in ly=y_sy th all 8 =0 m²_8m+16=0 ary equation Die 12) Given DE =) (6²_8D+ 16 )y Ito Ascili (marj²=0 m= 4,4 tet y= (a + Cqx) eth Aput y (o) plin → [4.4960)]2+0) pet 19=1 yll) at telet + cz cz[et+4et) -4C, et cz (set) so -44 - .4 ez de tous les ㅋ Y substi